The Warren Watch Company sells watches for $27, fixed costs are $100,000, and variable costs are $14 per watch.

  1. What is the firm's gain or loss at sales of 7,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
    $___

    What is the firm's gain or loss at sales of 20,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
    $___

  2. What is the break-even point (unit sales)? Round your answer to the nearest whole number.
    ___units

  3. What would happen to the break-even point if the selling price was raised to $32?
    ^^-SELECT AN ANSWER->>The result is that the break-even point is lower. OR The result is that the break-even point is higher. OR The result is that the break-even point remains unchanged.

  4. What would happen to the break-even point if the selling price was raised to $32 but variable costs rose to $27 a unit? Round your answer to the nearest whole number.
    ^^-SELECT AN ANSWER->>The result is that the break-even point increases. OR The result is that the break-even point decreases. OR The result is that the break-even point remains unchanged.
Answer #1

a.

Loss = (Selling price – Variable cost per unit) × Number of units – Fixed costs

            = (27 – 14) × 7,000 – 100,000

            = 13 × 7,000 – 100,000

            = 91,000 – 100,000

            = - 9,000 (Answer)

b

Gain = (Selling price – Variable cost per unit) × Number of units – Fixed costs

            = (27 – 14) × 20,000 – 100,000

            = 13 × 20,000 – 100,000

            = 260,000 – 100,000

            = 160,000 (Answer)

c.

BEP = FixedCost / (SellingPrice – Variable cost per unit)

            = 100,000 / (27 – 14)

            = 100,000 / 13

            = 7,692 (Answer)

d.

BEP = FixedCost / (SellingPrice – Variable cost per unit)

            = 100,000 / (32 – 14)

            = 100,000 / 18

            = 5,556 (Answer)

e.

Answer: BEP is lower.

It happens because 5,556 unit is lower than 7,692 unit.

f.

BEP = FixedCost / (SellingPrice – Variable cost per unit)

            = 100,000 / (32 – 27)

            = 100,000 / 5

            = 20,000 (Answer)

g.

Answer: BEP increases

It happens because 20,000 unit is higher than 7,692 unit.